OpenStudy (anonymous):

parallel lines on graph What is the solution to the system of equations represented by these two lines? A. (2, 3) B. (0, 4) C. (3, 2) D. (3, 0) @ganeshie8 @Directrix @SolomonZelman @UnkleRhaukus @uri @perl @Preetha @Jhannybean

OpenStudy (unklerhaukus):

the solution to the system, is the point of intersection

OpenStudy (anonymous):

is it c?

OpenStudy (unklerhaukus):

yer

OpenStudy (unklerhaukus):

alright

OpenStudy (unklerhaukus):

subsitute the second equation, into the first

OpenStudy (unklerhaukus):

y = -2x + 7 y = 1/3x 1/3x = -2x + 7

OpenStudy (anonymous):

first substitute the x/3=y of 2nd equation in first